Find the coordinates of R, Given S(-3,7) and the Midpoint of RS is (0,2.5)
Answer by lwsshak3(11628)

Find the coordinates of R, Given S(-3,7) and the Midpoint of RS is (0,2.5)
**
Midpoint formula beyween two points,(x1,y1) and (x2,y2): (x1+x2)/2, (y1+y2)/2
For given problem:
Between points R(x1,y1) and S(-3,7)
Midpoint: (x1-3)/2, (y1+7)/2=(0, 2.5)
(x1-3)/2=0
x1-3=0
x1=3
..
(y1+7)/2=2.5
y1+7=5
y1=-2
..
ans:
Coordinates of R: (3, -2)